The amount of hamburgers (X) and hotdogs (Y) in pounds that a man eats at a picnic is given by the following distribution:

f(x,y)=2x2+y2 for 0<x<1 and 0<y<1f(x,y)=2x2+y2 for 0<x<1 and 0<y<1

What is the expected amount of hamburger the man will eat?

Note x and y are squared
Answer #1

Oxx81, oxy< f(x,y) = 2x+y2 Hamburger cx) f(a) = fcx, 13 dy . - J ( x ) 44 f(x) = 2x + } ; 0<x</ E[x] = s refons dx = 1^(2x++)final answer = 2/3